# Exploring Local Memorization in Diffusion Models via Bright Ending Attention

### v1

@@ Abstract
In this paper, we identify and leverage a novel ‘bright ending’ (BE) anomaly in diffusion models prone to memorizing training images to address a new task: locating localized memorization regions within these models. BE refers to a distinct cross-attention pattern observed in text-to-image generations using diffusion models. Specifically, memorized image patches exhibit significantly greater attention to the end token during the final inference step compared to non-memorized patches. This attention map effectively highlights regions where the generated image replicates training data. Furthermore, driven by our observation that local memorization significantly underperforms in existing tasks of measuring, detecting, and mitigating memorization in diffusion models compared to global memorization, we propose a simple yet effective method to integrate BE and the results of the new localization task into these existing frameworks. This integration effectively improves their performances by narrowing the performance gap caused by local memorization. Our results not only demonstrate the successful execution of the new localization task but also establish new state-of-the-art performance across all existing tasks, underscoring the significance of the BE phenomenon.
@@ 1 Introduction
In response to these legal challenges, Carlini et al. (2023) and Somepalli et al. (2023a) proposed similarity metrics to evaluate memorization, and recent efforts (Somepalli et al., 2023b; Wen et al., 2024; Chen et al., 2024) have focused on developing strategies to detect and mitigate memorization, achieving notable success. However, these metrics and strategies adopt a global perspective, comparing entire generated images to training images. We identify a significant gap in these approaches when dealing with cases where only parts of the training image are memorized, which we refer to as local memorization, as opposed to cases where the entire training image is memorized, termed global memorization. Specifically, current methods can be easily tricked in local memorization scenarios by introducing diversity in non-memorized regions, resulting in many false negatives during evaluation and detection. Consequently, this failure to correctly identify local memorization may prevent the activation of mitigation strategies. Motivated by this research gap, we developed a localization insight as a novel view of the memorization issues, where we argue that improved metrics and strategies should concentrate exclusively on the locally memorized regions, as these areas pose risks. In contrast, unmemorized parts of the image, which do not present risks, should be completely ignored.
@@
Building on this localization insight, we introduce a new task: extracting localized memorized regions. This task is crucial for better understanding local memorization and developing targeted strategies to reduce the associated performance gap. To accomplish this task effectively and efficiently, we introduce the concept of the ‘bright ending’ (BE), a phenomenon observed in the cross-attention maps of text-to-image diffusion models. The bright ending occurs when the end token in the final inference step exhibits abnormally high attention scores on specific image patches in memorized generations compared to non-memorized ones. This distinctive pattern effectively highlights the regions where the model has memorized the training data, requiring only a single inference pass and without relying on access to the training data. To our knowledge, this is the first method to achieve such precise and efficient localization, emphasizing the significance of BE. Finally, we propose a simple yet effective method to integrate BE and its extracted local memorization masks into existing state-of-the-art memorization evaluation, detection, and mitigation strategies. Our extensive experiments demonstrate that this integration significantly enhances the performance of these existing tasks by narrowing the gap caused by local memorization, further underscoring the contribution of BE.
@@
In summary, our contributions are three-fold: 1) We analyze the performance gap between local and global memorization in existing tasks and introduce a novel localized view of memorization in diffusion models. 2) We pioneer a new task of locating localized memorization regions and propose the novel ‘bright ending’ (BE) phenomenon, which is the first method to successfully achieve this task using only a single inference process without requiring access to training data and conducting a nearest neighbor search over all ground truth images. 3) By integrating the local memorization mask extracted using BE into existing evaluation, detection, and mitigation strategies, we refine these approaches to focus on local perspectives, in contrast to the global focus of all existing strategies. This refinement narrows the gap caused by local memorization and sets a new state-of-the-art in these existing tasks. This simple and effective incorporation further demonstrates the power of BE.
@@
Figure 1: Memorization in diffusion models can occur both globally (right) and locally (left). The first row displays the memorized training image, the second row shows the corresponding generated image, and the third row highlights the local memorization mask extracted using ‘bright ending’.
@@ 2.1 Diffusion Models
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})$ (1)
@@
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$ (3)
@@
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t}),\sigma_{\theta}^{2}(x_{t})\mathbf{I})$ (4)
$p_{\theta}(x_{T:0})=p(x_{T})\prod_{t=T}^{1}p_{\theta}(x_{t-1}|x_{t})$ (5)
@@
$\mathcal{L}=\mathbb{E}_{t\in[1,T],\epsilon\sim\mathcal{N}(0,\mathbf{I})}[\left\|\epsilon_{t}-\epsilon_{\theta}(x_{t},t)\right\|_{2}^{2}]$ (6)
@@ 3 A localized view of memorization in diffusion models
Figure 2: Left: Variations in non-memorized regions significantly impact similarity scores with existing global metrics. Right: Local memorization is more challenging to detect than global memorization, with more failure cases observed using the current state-of-the-art detection method.
@@
We observe that all existing studies on memorization in diffusion models are from a global perspective, where the effectiveness of any proposed detection methods and mitigation strategies are evaluated using global metrics, such as L2 distance or Self-Supervised Copy Detection (SSCD) (Pizzi et al., 2022). These metrics measure similarity between entire generated images and training images rather than focusing on locally memorized areas. However, these metrics are not robust to variations outside the locally memorized areas and can be easily tricked. A generated image may memorize only a local area of the training image while making the remaining areas novel and diverse, leading to false negatives. Fig. 2 (left) illustrates failures of the global metric approach, where variations in non-memorized regions significantly reduce the similarity scores measured by SSCD. The current standard considers SSCD scores greater than 0.5 as memorization. This results in some images being falsely identified as non-memorized despite all being locally memorized, with differences only in the non-memorized regions.
@@
Current detection and mitigation strategies are evaluated using global metrics and are thus developed from a global perspective. For instance, Wen et al. (2024) distinguishes memorized cases from unmemorized ones by utilizing the magnitude over the global latent space (Eq. 7), with memorized cases exhibiting larger magnitudes. This detection method has outperformed its predecessors, setting the current state-of-the-art. However, this approach is less effective for local memorization cases than for global ones, as variations in non-memorized regions impact magnitude computations, increasing variance and diminishing the reliability of the detection signal, which leads to more false negatives. To verify this theory, we visualized the density plot of magnitudes for all three scenarios (global, local, and no memorization) in Fig. 2 (right) and Fig. 3 and observed that local cases indeed have smaller magnitudes than global ones, making them closer in magnitude to unmemorized cases and harder to detect. Regarding mitigation, local memorization also underperforms. Wen et al. (2024) employs prompt engineering when a positive detection is made, optimizing the prompt using global magnitude as the loss function, which is, therefore, also sensitive to variations in non-memorized regions. Additionally, the compromised global detection method, which produces more false negatives, often leaves the mitigation strategy untriggered, increasing the risk of unaddressed memorization and potentially leading to legal challenges.
$d=\frac{1}{T}\sum_{t=1}^{T}\left\|\left(\varepsilon_{\theta}(x_{t},e_{p})-\varepsilon_{\theta}(x_{t},e_{\phi})\right)\right\|_{2}\vspace{-0.2cm}$ (7)
@@ 4 Bright ending
The observed performance gap caused by local memorization has prompted us to explore a new task: extracting localized memorized regions as a mask. This task is essential for investigating memorization from a local perspective, and the resulting mask can be integrated into existing evaluation, detection, and mitigation strategies. This integration refines these strategies into localized approaches, effectively reducing the performance gap associated with local memorization. One naive method to create such a local memorization mask is to compare generated images with the closest training image. However, this reliance on training data presents limitations: (1) It compromises privacy for a task aimed at preventing the memorization of training data and assumes that the user of the pre-trained model has access to the training data; (2) It requires significantly more computation due to the need for nearest neighbor searches on the large LAION training dataset. Consequently, this naive approach diminishes its practical significance when the mask is used in detection and mitigation algorithms. Therefore, we aim to extract such masks without using training data, relying solely on the pre-trained model’s memory to identify distinguishing memorization patterns.
Recent work has highlighted the important role of prompts in causing memorization. For example, Somepalli et al. (2023b) identified that duplicated prompts during the training of text-to-image Stable Diffusion models make the prompts become the “keys” to the models’ memory. Similarly, Wen et al. (2024) observed that prompts prone to memorization tend to result in larger text-conditional noise predictions during inference than non-memorization cases. Additionally, Chen et al. (2024) successfully utilized classifier-free guidance to steer generations away from memorized prompts, improving mitigation. Inspired by these findings, we focused on using prompts to extract our desired local mask. This naturally led us to explore the cross-attention mechanism in the U-Net architecture within diffusion models, which links prompts to specific attention-focusing areas in the generated images. One of our key intuitions is that the pre-trained diffusion model’s memories resulting from overfitting should differ from those learned through generalization. Memorized generations are inflexible and follow a fixed denoising trajectory that is observably different from non-memorized scenarios. This is evidenced by Wen et al. (2024), who observed that in memorized cases, the magnitude of text-conditional noise prediction during the denoising trajectory is abnormally higher than in non-memorization cases.
@@
Figure 5: Left: Box plot showing the distribution of end token cross-attention scores during the final inference step for different memorization types. The plot is based on 9,600 generated images, with 16 images generated for 300 memorized and 300 non-memorized prompts each. It clearly distinguishes the three types: non-memorized cases have attention scores close to zero, while memorized cases exhibit abnormally high scores. Global memorization shows higher scores than local memorization due to the larger memorized regions. This further validates the ‘bright ending’ observation. Right: Violin plots displaying the impact on magnitudes for different memorization types.
@@
Inspired by our intuition, we visualized each token’s cross-attention maps in memorized and non-memorized prompts to identify distinguishing patterns (Fig. 4). We discovered an interesting phenomenon in pre-trained text-to-image Stable Diffusion models: during the final denoising step, the end token typically shows a dark cross-attention map. Specifically, these dark maps often contain only extremely low-attention finer edges or random scatter points. Since the end token summarizes all semantic meanings in the prompt, this indicates that when the noised image is very close to a clean image, the denoiser’s attention on text-conditioning shifts from focusing on its semantic meanings to concentrating only on fine details.
However, when the model exhibits memorization, we observe an anomaly where the end token displays abnormally high cross-attention scores at the end of denoising steps — a phenomenon we refer to as the ‘bright ending’ (BE). Specifically, it focuses on the coarser structure of image patches rather than finer details or random scatter points. This indicates that the denoiser remains focused on text-conditioning even during the final inference step, demonstrating overfitting to the text-conditioning. This aligns with the finding in Somepalli et al. (2023b) that memorized generations result from overfitting to duplicated caption-conditioning during training, thus following a fixed noise trajectory during inference. Thus, BE is a robust indicator for differentiating memorized generations from unmemorized ones, with bright regions identifying the memorized areas. This effectively functions as an automatic memorization mask, acting as a global mask during global memorization and a local mask during local memorization, as can be visually observed in Fig. 4 and 1. To further validate this observation, we used 300 memorized prompts and 300 non-memorized prompts to generate 16 images for each prompt, resulting in a total of 9600 generations. We then employed a box plot (Fig. 5) to visualize the distribution of the attention scores of the end token during the final inference step for these generated images. We observed a clear distinction among the three memorization types: non-memorized cases consistently had attention scores close to zero, while memorized cases exhibited abnormally high scores. Furthermore, global memorization showed higher scores than local memorization due to the larger regions being memorized in global cases compared to local ones.

### v2

@@ Abstract
Text-to-image diffusion models have achieved unprecedented proficiency in generating realistic images. However, their inherent tendency to memorize and replicate training data during inference raises significant concerns, including potential copyright infringement. In response, various methods have been proposed to evaluate, detect, and mitigate memorization. Our analysis reveals that existing approaches significantly underperform in handling local memorization, where only specific image regions are memorized, compared to global memorization, where the entire image is replicated. Also, they cannot locate the local memorization regions, making it hard to investigate locally. To address these, we identify a novel “bright ending” (BE) anomaly in diffusion models prone to memorizing training images. BE refers to a distinct cross-attention pattern observed in text-to-image diffusion models, where memorized image patches exhibit significantly greater attention to the final text token during the last inference step than non-memorized patches. This pattern highlights regions where the generated image replicates training data and enables efficient localization of memorized regions. Equipped with this, we propose a simple yet effective method to integrate BE into existing frameworks, significantly improving their performance by narrowing the performance gap caused by local memorization. Our results not only validate the successful execution of the new localization task but also establish new state-of-the-art performance across all existing tasks, underscoring the significance of the BE phenomenon.
@@ 1 Introduction
In response to these legal challenges, Carlini et al. (2023) and Somepalli et al. (2023a) proposed similarity metrics to evaluate memorization, and recent efforts (Somepalli et al., 2023b; Wen et al., 2024; Chen et al., 2024) have focused on developing strategies to detect and mitigate memorization, achieving notable success. However, these metrics and strategies adopt a global perspective, comparing entire generated images to training images. We identify a significant gap in these approaches when dealing with cases where only parts of the training image are memorized, which we refer to as local memorization, as opposed to cases where the entire training image is memorized, termed global memorization. Specifically, current methods can be easily tricked in local memorization scenarios by introducing diversity in non-memorized regions, resulting in many false negatives during evaluation and detection. Consequently, this failure to correctly identify local memorization may prevent the activation of mitigation strategies. Motivated by this research gap, we propose a novel view of the memorization issues, where we argue that improved metrics and strategies should concentrate exclusively on the locally memorized regions, as these areas pose risks. In contrast, unmemorized parts of the image, which do not present risks, should be completely ignored.
@@
Building on this localization insight, we introduce a new task: extracting localized memorized regions. This task is crucial for better understanding local memorization and developing targeted strategies to reduce the associated performance gap. To accomplish this task effectively and efficiently, we introduce the concept of the ‘bright ending’ (BE), a phenomenon observed in the cross-attention maps of text-to-image diffusion models. The bright ending occurs when the end text token in the final inference step exhibits abnormally high attention scores on specific image patches in memorized generations compared to non-memorized ones. This distinctive pattern effectively highlights the regions where the model has memorized the training data, requiring only a single inference pass without relying on access to the training data. To our knowledge, this is the first method to achieve such precise and efficient localization, emphasizing the significance of BE. Finally, we propose a simple yet effective method to integrate BE and its extracted local memorization masks into existing state-of-the-art memorization evaluation, detection, and mitigation strategies. Extensive experiments demonstrate that this integration significantly enhances the performance of these existing tasks by narrowing the gap caused by local memorization, further underscoring the contribution of BE.
@@
In summary, our contributions are three-fold: 1) We analyze the performance gap between local and global memorization in existing tasks and introduce a novel localized view of memorization in diffusion models. 2) We propose a new task of detecting localized memorization regions and propose the novel ‘bright ending’ (BE) phenomenon, which is the first method to successfully solve this task using only a single inference process without requiring access to training data and conducting a nearest neighbor search over all ground truth images. 3) By integrating the local memorization mask extracted using BE into existing evaluation, detection, and mitigation strategies, we refine these approaches to focus on local perspectives, in contrast to the global focus of all existing strategies. This refinement narrows the gap caused by local memorization and sets a new state-of-the-art in these existing tasks. This simple and effective incorporation further demonstrates the power of BE.
@@
Figure 1: Memorization in diffusion models can occur both globally (right) and locally (left). The first row displays the memorized training images, the second row shows the corresponding generated images, and the third row shows the local memorization masks extracted using ‘bright ending’, where brightness values quantify the memorization effect. Most pixel values in masks on the right are brighter due to global memorization compared to masks on the left due to local memorization.
@@ 2.1 Diffusion Models
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I}),$ (1)
@@
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I}),$ (3)
@@
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t}),\sigma_{\theta}^{2}(x_{t})\mathbf{I}),$ (4)
$p_{\theta}(x_{T:0})=p(x_{T})\prod_{t=T}^{1}p_{\theta}(x_{t-1}|x_{t}),$ (5)
@@
$\mathcal{L}=\mathbb{E}_{t\in[1,T],\epsilon\sim\mathcal{N}(0,\mathbf{I})}[\left\|\epsilon_{t}-\epsilon_{\theta}(x_{t},t)\right\|_{2}^{2}],$ (6)
@@ 3 A localized view of memorization in diffusion models
Figure 2: For local memorization cases, variations in non-memorized regions significantly impact existing global similarity and detection measures. Left: Using the current similarity metric SSCD, the second row shows memorized images with similarity scores above 0.5, and the third row shows non-memorized images with scores below 0.5. However, the difference between the second and third rows is minimal, differing only in a local portion, and are, in fact, both local memorization cases, which demonstrates the failure of the global SSCD metric. Right: Currently, magnitude is used to detect memorization, where a higher magnitude corresponds to a greater memorization risk. The distribution reveals a noticeable performance gap: local memorization cases are more challenging to detect than global ones, as their magnitude values overlap more with those of non-memorized cases.
@@
We observe that all existing studies on memorization in diffusion models are from a global perspective, where the effectiveness of any proposed detection methods and mitigation strategies are evaluated using global metrics, such as the L2 distance or SSCD scores (i.e., cosine similarity of embeddings extracted using the Self-Supervised Copy Detection (SSCD) method (Pizzi et al., 2022)). These metrics measure similarity between entire generated and training images rather than focusing on locally memorized areas. However, these metrics are not robust to variations outside the locally memorized areas and can be easily tricked. Specifically, variations in non-memorized regions can significantly impact these similarity scores despite the similar local memorization regions. A generated image may memorize only a local area of the training image while making the remaining areas novel and diverse, leading to false negatives. Fig. 2 (left) illustrates failures of the global metric approach, where variations in non-memorized regions significantly reduce the similarity scores measured by SSCD. The current standard considers SSCD scores greater than 0.5 as memorization. This results in some images being falsely identified as non-memorized despite all being locally memorized, with differences only in the non-memorized regions.
@@
Current detection and mitigation strategies are evaluated using global metrics and are thus developed from a global perspective. For instance, Wen et al. (2024) distinguishes memorized cases from unmemorized ones by utilizing the magnitude over the global latent space (Eq. 7):
$D=\frac{1}{T}\sum_{t=1}^{T}\left\|\left(\varepsilon_{\theta}(x_{t},e_{p})-\varepsilon_{\theta}(x_{t},e_{\phi})\right)\right\|_{2},$ (7)
where $e_{p}$ represents the embedding of the user-provided prompt, $e_{\phi}$ denotes the embedding of an empty string, $\varepsilon_{\theta}(x_{t},\cdot)$ is the conditional latent-space noise prediction at time $t$, and $D$ is the computed magnitude. The detection method identifies memorized cases by their larger magnitudes $D$ compared to non-memorized cases. This approach surpasses previous methods, establishing the current state-of-the-art. However, this approach is less practical for local memorization cases than for global ones, as variations in non-memorized regions impact magnitude computations, increasing variance and diminishing the reliability of the detection signal, which leads to more false negatives. To verify this theory, we visualize the density plot of magnitudes for all three scenarios (global, local, and no memorization) in Fig. 2 (right) and Fig. 3 and observed that local cases indeed have smaller magnitudes than global ones, making them closer in magnitude to unmemorized cases and harder to detect. Regarding mitigation, local memorization also underperforms. Wen et al. (2024) employs prompt engineering upon a positive detection, i.e., when $D$ exceeds a threshold $\lambda$, the generation is identified as memorized. Subsequently, $e_{p}$ is optimized via gradient descent until $D\leq\lambda$, using the global magnitude $D$ as the loss function. Consequently, this optimization process is also sensitive to variations in non-memorized regions. Additionally, the compromised global detection method, which produces more false negatives, often leaves the mitigation strategy untriggered, increasing the risk of unaddressed memorization and potentially leading to legal challenges.
@@ 4 Bright ending
The observed performance gap caused by local memorization has prompted us to explore a new task: extracting localized memorized regions as a mask. This task is essential for investigating memorization from a local perspective, and the resulting mask can be integrated into existing evaluation, detection, and mitigation strategies. This integration refines these strategies into localized approaches, effectively reducing the performance gap associated with local memorization. One naive method to create such a local memorization mask is to compare generated images with the closest training image. However, this reliance on training data has following limitations: (1) It compromises privacy for a task aimed at preventing the memorization of training data and assumes that the user of the pre-trained model has access to the training data; (2) It requires significantly more computation due to the need for nearest neighbor searches on the large LAION training dataset. Consequently, this naive approach diminishes its practical significance when the mask is used in detection and mitigation algorithms. Therefore, we aim to extract such masks without using training data, relying solely on the pre-trained model’s memory to identify distinguishing memorization patterns.
Recent work has highlighted the important role of prompts in causing memorization. For example, Somepalli et al. (2023b) identified that duplicated prompts during the training of text-to-image Stable Diffusion models make the prompts become the “keys” to the models’ memory. In other words, when a model encounters a frequently seen prompt, it retrieves and generates images that closely resemble memorized training examples, much like a key unlocking stored content. Similarly, Wen et al. (2024) observed that prompts prone to memorization tend to result in larger text-conditional noise predictions during inference than non-memorization cases. Additionally, Chen et al. (2024) successfully utilized classifier-free guidance to steer generations away from memorized prompts, improving mitigation. Inspired by these findings, we focus on using prompts to extract our desired local mask. This naturally led us to explore the cross-attention mechanism in the U-Net architecture within diffusion models, which links prompts to specific attention-focusing areas in the generated images. One of our key intuitions is that the pre-trained diffusion model’s memories resulting from overfitting should differ from those learned through generalization. Memorized generations are inflexible and follow a fixed denoising trajectory that is observably different from non-memorized scenarios. This is evidenced by Wen et al. (2024), who observed that in memorized cases, the magnitude of text-conditional noise prediction during the denoising trajectory is abnormally higher than in non-memorization cases.
@@
Figure 5: Left: Box plot showing the distribution of end token cross-attention scores during the final inference step for different memorization types. The plot is based on 9,600 generated images, with 16 images generated for 300 memorized and 300 non-memorized prompts each. It clearly distinguishes the three types: non-memorized cases have attention scores close to zero, while memorized cases exhibit abnormally high scores. Global memorization shows higher scores than local memorization due to the larger memorized regions. This further validates the ‘bright ending’ observation. Right: Violin plots displaying the impact on magnitudes (detection signal) for different memorization types.
@@
Inspired by our intuition, we visualize each token’s cross-attention maps in memorized and non-memorized prompts to identify distinguishing patterns (Fig. 4). We discover an interesting phenomenon in pre-trained text-to-image Stable Diffusion models: during the final denoising step, the end token typically shows a dark cross-attention map. Specifically, these dark maps often contain only extremely low-attention finer edges or random scatter points. Since the end token summarizes all semantic meanings in the prompt, this indicates that when the noised image is very close to a clean image, the denoiser’s attention on text-conditioning shifts from focusing on its semantic meanings to concentrating only on fine details.
However, when the model exhibits memorization, we observe an anomaly where the end token displays abnormally high cross-attention scores at the end of denoising steps — a phenomenon we refer to as the ‘bright ending’ (BE). Specifically, it focuses on the coarser structure of image patches rather than finer details or random scatter points. This indicates that the denoiser remains focused on text-conditioning even during the final inference step, demonstrating overfitting to the text-conditioning. This aligns with the finding in Somepalli et al. (2023b) that memorized generations result from overfitting to repeated captions used as conditioning signals for text-to-image generation during training, thereby following a fixed noise trajectory during inference. Thus, BE is a robust indicator for differentiating memorized generations from unmemorized ones, with bright regions identifying the memorized areas. This effectively functions the BE attention map as an automatic memorization mask, acting as a global mask during global memorization and a local mask during local memorization, as can be visually observed in Fig. 4 and 1. To further validate this observation, we used 300 memorized prompts from Webster (2023) and 300 non-memorized prompts to generate 16 images for each prompt, resulting in a total of 9600 generations. We then employ a box plot (Fig. 5) to visualize the distribution of the attention scores of the end token during the final inference step for these generated images. We observe a clear distinction among the three memorization types: non-memorized cases consistently had attention scores close to zero, while memorized cases exhibited abnormally high scores. Furthermore, global memorization showed higher scores than local memorization due to the larger regions being memorized in global cases compared to local ones.